write an equation of the line containing the given point and perpendicular to the given line.
(5,-6); 8x+9y=7

Answer by mananth(16949)

write an equation of the line containing the given point and perpendicular to the given line.
(5,-6); 8x+9y=7
...
8x+9y=7
9y=-8x+7
/9
y=(-8x/9)+7/9
slope = -8/9
...
slope of line perpendicular to above line will be 9/8
..
m=9/8,(5,-6)
Find b in y=mx +b
plug values
-6=9/8*5+b
-6=45/8+b
b=-6-45/8
b=-93/8
..
y=(9x/8)-(93/8)
...
m.ananth@hotmail.ca